Solve r and interpret the result


X 2,4,6,7,10


Y 8,10,12,6,16



In the given table on the below, solve for Pearson r and interpret the result


X 80,84,86,87,89,90,91,93,94,96


Y 78,83,80,84,89,90,88,91,93,96

1
Expert's answer
2022-06-13T18:05:24-0400

a) In order to compute the regression coefficients, the following table needs to be used:


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c}
 & X & Y & XY & X^2 & Y^2 \\ \hline
 & 2 & 8 & 16 & 4 & 64 \\
 \hdashline
 & 4 & 10 & 40 & 16 & 100 \\
 \hdashline
 & 6 & 12 & 72 & 36 & 144 \\
 \hdashline
 & 7 & 6 & 42 & 49 & 36 \\
 \hdashline
 & 10 & 16 & 160 & 100 & 256 \\
 \hdashline
Sum= & 29 & 52 & 330 & 205 & 600 \\
 \hdashline
\end{array}"




"\bar{X}=\dfrac{1}{n}\sum _{i}X_i=\dfrac{29}{5}=5.8"




"\bar{Y}=\dfrac{1}{n}\sum _{i}Y_i=\dfrac{52}{5}=10.4"




"SS_{XX}=\sum_iX_i^2-\dfrac{1}{n}(\sum _{i}X_i)^2""=205-\dfrac{29^2}{5}=36.8"




"SS_{YY}=\sum_iY_i^2-\dfrac{1}{n}(\sum _{i}Y_i)^2""=600-\dfrac{(52)^2}{5}=59.2"




"SS_{XY}=\sum_iX_iY_i-\dfrac{1}{n}(\sum _{i}X_i)(\sum _{i}Y_i)""=330-\dfrac{29(52)}{5}=28.4"




"r=\dfrac{SS_{XY}}{\sqrt{SS_{XX}SS_{YY}}}=\dfrac{28.4}{\sqrt{36.8(59.2)}}"




"=0.608462"




"0.4<r<0.7"

Moderate positive correlation



"m=slope=\dfrac{SS_{XY}}{SS_{XX}}=\dfrac{28.4}{36.8}=0.7717""n=\bar{Y}-m\bar{X}=10.4-\dfrac{28.4}{36.8}(5.8)=5.9239"


The regression equation is:


"y=5.9239+0.7717x"



(b) In order to compute the regression coefficients, the following table needs to be used:


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c}
 & X & Y & XY & X^2 & Y^2 \\ \hline
 & 80 & 78 & 6240 & 6400 & 6084 \\
 \hdashline
 & 84 & 83 & 6972 & 7056 & 6889 \\
 \hdashline
 & 86 & 80 & 6880 & 7396 & 6400 \\
 \hdashline
 & 87 & 84 & 7308 & 7569 & 7056 \\
 \hdashline
 & 89 & 89 & 7921 & 7921 & 7921 \\
 \hdashline
 & 90 & 90 & 8100 & 8100 & 8100 \\
 \hdashline
 & 91 & 88 & 8008 & 8281 & 7744 \\
 \hdashline
 & 93 & 91 & 8463 & 8649 & 8281 \\
 \hdashline
 & 94 & 93 & 8742 & 8836 & 8649 \\
 \hdashline
 & 96 & 96& 9216 & 9216 & 9216 \\
 \hdashline
Sum= & 890 & 872 & 77850 & 79424 & 76340 \\
\end{array}"




"\bar{X}=\dfrac{1}{n}\sum _{i}X_i=\dfrac{890}{10}=89"




"\bar{Y}=\dfrac{1}{n}\sum _{i}Y_i=\dfrac{872}{10}=87.2"




"SS_{XX}=\sum_iX_i^2-\dfrac{1}{n}(\sum _{i}X_i)^2""=79424-\dfrac{890^2}{10}=214"




"SS_{YY}=\sum_iY_i^2-\dfrac{1}{n}(\sum _{i}Y_i)^2""=76340-\dfrac{(872)^2}{10}=301.6"




"SS_{XY}=\sum_iX_iY_i-\dfrac{1}{n}(\sum _{i}X_i)(\sum _{i}Y_i)""=77850-\dfrac{890(872)}{10}=242"




"r=\dfrac{SS_{XY}}{\sqrt{SS_{XX}SS_{YY}}}=\dfrac{242}{\sqrt{214(301.6)}}"




"=0.952561"


Strong positive correlation